Words
102
Reading
1 min
7y
Ok I am doing some phone math with me so bare with me. Let’s say only 1 witness is running this code.
There is a 1/21 chance they are in the last position and a 1/21 chance they are in the first position in the next round. 1/21 * 1/21 = 0.00226757 probability of this happening.
123 second per 2 complete rounds / 0.00226757 = 54243 second before this occurs = 15.068 hours.
So we could expect this transaction to go through in less than one day. Interesting. This would obviously be faster if more than 1 witness are running this code.
RE: Theory: How to side-step soft-consensus